# Date of publication: December 8th, 2025

### v1

@@ Abstract
Mountain huts are buildings located at high altitude, offering a place for hikers and providing shelter. Energy supply on mountain huts is still an open issue. Using renewable energies could be an appropriate solution. Jiujiu Cabins, a famous mountain hut in Shei-Pa National Park, Taiwan, has operated an off-grid solar energy storage system (ESS) with lead-acid batteries. In 2021, a serious system failures took place, leading to no electricity. After an detailed on-site survey, a reorganization and repair project implemented, the energy system came back to operate normally. Meanwhile, a eco-friendly lithium iron phosphate battery (LFP battery) ESS replaces part of the lead-acid battery ESS, forming a hybrid ESS, making a better and green off-grid solar ESS. In this case report, the energy architecture, detailed descriptions, and historical status of the system are provided. An on-site survey of the failed energy system, a system improvement project, and future plan are listed.
@@
microgrid, renewable energy, photovoltaic system, energy storage system, hybrid energy storage system, lithium-ion battery, lithium iron phosphate battery, high mountain, engineering.
@@ I Introduction
Mountain huts are buildings located at high altitude, in mountainous terrain, generally accessible only by foot, intended to provide food and shelter to mountaineers, hikers, and climbers [1]. They are known by many names, such as alpine hut, mountain shelter, mountain refuge, mountain lodge, mountain hostel, and mountain cabin. Mountain huts can provide some functionalities and services. The most important function is to provide shelter and simple sleeping berths for hikers to overcome the cold mountain environment at night. Some mountain huts are not staffed, particularly in remote areas. However, others have staff who provide hot meals, drinks, and other services, such as activity spaces (for holding lectures, etc.), clothing sales, and small items.
Energy supply in high mountains remains an open issue to be solved. Generally, grid connection is impractical, because the establishment and maintenance costs are too high. It also causes many environmental problems, e.g., the destruction of the natural ecology in the mountain. Diesel generators as energy suppliers have been used in high mountains for a long time despite their air pollution and noise problems. Due to the development of renewable energy (such as solar, wind, and hydropower), the usage of diesel generators is reduced, lowering the emissions of greenhouse gases (GHGs). However, owing to their fluctuating nature, most renewable energy sources exhibit intermittent features. To deal with the problem of renewable energy intermittency, energy storage systems (ESSs) are necessary [2, 3]. In the past, lead-acid batteries were heavily used as ESSs, accompanied by many environmental issues, e.g., poisoning, leaks, contamination of the environment, and damage to the ecosystem [4, 5]. Recently, lithium-ion (Li-ion) batteries [6, 7], as a greener alternative, have started to replace lead-acid batteries in ESSs.
@@
In this manuscript, a brief introduction of the Jiujiu Cabins is given in Sec. II. In Sec. III, the current status of the energy architecture of the off-grid solar energy storage system is presented with a detailed description of each component. A historical development of energy systems is listed in Sec. IV. In 2021, Jiujiu Cabins experienced a brief period of power outage, and the investigation results of the system failure are listed in Sec. V. How to overhaul and improve the system, and the engineering work to restore power, is covered in Sec. VI. Possible future system upgrade plans are listed in Sec. VII. Summary and Outlook are given at last (Sec. VIII).
@@ II-A Current status
Jiujiu Cabins (coordinate: $24^{\circ}28^{\prime}08.7^{\prime\prime}$N, $121^{\circ}12^{\prime}29.9^{\prime\prime}$E) is an accommodation mountain lodge in Taiwan, located in Shei-Pa National Park, in Tai’an Township, Miaoli County, at an altitude of 2,699 m (8,855 ft). It is named after the number 99 in its altitude [8]. The lodge complex belongs to the Hsinchu Branch, Forestry and Nature Conservation Agency, Ministry of Agriculture. It includes Chengkung Yurts, a restaurant (reception room), Longmen Inns, a machine room, toilet, etc., all of which are separate buildings (as shown in Fig. 2). Jiujiu Cabins uses a photovoltaic system and is equipped with a lithium-ion battery energy storage system to cope with the intermittent nature of renewable energy [8]. It can currently accommodate 150 people [9, 10]. It has clean running water (but currently no hot water) and a mobile phone communication point [11].
@@
Figure 2: Aerial photography of Jiujiu Cabins (2021). The lodge complex. It includes Chengkung Yurts, a restaurant (reception room), Longmen Inns, a machine room, toilet, etc., all of which are separate buildings. (Photographer: Yu-Chun Lin)
@@ III Architecture of current energy system
Due to the difficulty of power grid connection in high-altitude areas, energy supply remains a challenging problem. Off-grid solar energy storage systems [18, 19] provide an appropriate solution for using renewable energy in high-altitude areas and have been proven [18].
In order to restore the operation of the energy system, the Hsinchu Forest District Office, Forestry Bureau (now the Hsinchu Branch, Forestry and Nature Conservation Agency, Ministry of Agriculture, Taiwan) commissioned Super Double Power Technology Co., Ltd. to improve the existing solar energy system of Jiujiu Cabins in 2021 [20]. After restructuring and improvement, the off-grid solar energy storage system of Jiujiu Cabins is composed of two independent power supply systems combined with the same energy storage system architecture, as shown in Fig. 3. Among them, the first PV array inputs solar energy to hybrid solar inverters, which can transmit power to alternating current (AC) loads. The second PV array can transmit power to another set of AC loads through another hybrid solar inverter. The two sets of hybrid solar inverters share the same hybrid ESS, which consists of a newly established LFP battery [21] ESS and an existing lead-acid battery ESS. In addition, a new backup power input terminal is added, which can be connected to an external generator as a backup power source.
@@ III-A Photovoltaic systems (PV systems)
In Jiujiu Cabins, there are two PV arrays, providing a total power generation capacity of 5.3 kWp. The first channel of the PV array is a ground-mounted PV system with a power generation capacity of 2.1 kWp. It is composed of 28 75Wp solar modules (SIEMENS Solar module SP75), as shown in Fig. 4. The second channel of the PV array is a ground-mounted PV system with a power generation capacity of 3.2 kWp. It is composed of 40 80Wp solar modules (SHARP Solar Module NE-80E2E), as shown in Fig. 5.
Figure 4: Ground-mounted PV system 1 PV array 1 is a ground-mounted PV system with a power generation capacity of 2.1 kWp. It is composed of 28 75Wp solar modules.
Figure 5: Ground-mounted PV system 2 PV array 2 is a ground-mounted PV system with a power generation capacity of 3.2 kWp. It is composed of 40 80Wp solar modules.
@@ III-F Backup power system
A diesel generator could be the backup power system. When the power from the PV system is insufficient (e.g., rainy days or cloudy days), the diesel generator can supply the insufficient power to the lodge. When the system encounters failures (e.g., temporary shutdown of the hybrid solar inverter), the diesel generator can supply the full power to the lodge.
@@ IV Historical development of energy systems
The Jiujiu Cabins had two existing electrical systems, designated as Old System 1 and Old System 2. These systems were likely installed by different manufacturers at different times. Due to a malfunction in the Old System 1, power was unavailable. Another manufacturer then installed Old System 2 to supply power to the cabins’ 110 VAC loads. However, the inverter in Old System 2 also failed due to aging, causing power supply anomalies and preventing it from outputting electricity.
In order to restore the operation of the energy system, the Hsinchu Forest District Office commissioned Super Double Power Technology Co., Ltd. to reorganize, repair, and improve the existing solar energy system of Jiujiu Cabins in 2021 [20]. After the improvement project, it’s the current energy architecture, which is described in Sec. III in detail.
@@
Before Total: 5.3 kWp (PV) 15 kWh (Lead-acid) 3 kW Yes
@@ V On-site survey of energy systems
The Old System 1 can be considered as a microgrid system built on a 48 VDC DC bus. Solar panels (2.1 kWp) generate electricity, which is then charged via a solar charger (PWM) to a 48 VDC lead-acid battery energy storage system. A 48 VDC inverter converts the DC power to 110 VAC for output. Based on the property label, its installation date is likely Jan. 20, 2000.
The Old System 2 can be considered as a microgrid system built on a 24 VDC DC bus. After generating electricity from the solar panels (3.2 kWp), the electricity is charged to the 24 VDC lead-acid battery energy storage system through a solar charger (PWM). The DC power is then converted to 110 VAC AC power output by a 24 VDC inverter, with a rated power of 3 kW.
@@ VI-A Key issues
    The two existing PV arrays can be reused, but the series and parallel connection needs to be adjusted within the inverter specifications.
@@ VI-B Main engineering subprojects
The solution involves restructuring the solar power lines, upgrading the inverters, restructuring the power supply lines at the cabins, and increasing the capacity of the energy storage system. However, this is not a simple matter of adding more batteries in parallel. It involves the parallel connection of old and new ESSs and the integration of the entire energy system.
The main engineering subprojects are: (1) adding LFP battery modules, (2) updating inverters, (3) modifying and reorganizing the existing PV array 1, (4) modifying and reorganizing the existing PV array 2, (5) changing control boxes (changing AC/DC lines), and (6) materials (hardware parts/connectors/DC lines).
The following describes each subproject in detail:
@@
    Modifying and reorganizing the existing PV array 1. The PV series and parallel configuration of PV array 1 (2.1 kWp) was modified, so that its operating voltage is in the range of 90 230 VDC and its open circuit voltage (OCV) is less than 250 VDC, which can meet the inverter input specifications.
@@
    Modifying and reorganizing the existing PV array 2. The PV series and parallel configuration of PV array 2 (3.2 kWp) was modified, so that its operating voltage is in the range of 90 230 VDC and its OCV is less than 250 VDC, which can meet the inverter input specifications.
@@
    Materials (hardware parts/connectors/DC lines). (a) The old hardware parts/connectors/DC lines were removed; (b) hardware parts/connectors/DC lines were added according to the current decoration situation, including DC surge absorbers, DC dedicated leakage circuit breakers, coulomb meters, cloud EMS, and generator charging input terminals.
@@ VI-C EMS data analysis
In this project, due to budget constraints, the self-developed cloud-native energy management system only records the total voltage, input/output current, and input/output power of the energy storage system. Because the cabins’ energy usage has a certain cyclical characteristic on a daily basis, a one-day data analysis is used to assess the system’s condition, as shown in Fig. 8.
@@
The 24-hour data shows that the highest total voltage of the ESS does not exceed 54 VDC (as shown in Fig. 8(a)), indicating that the BMS provides stable control over the LFP battery, keeping it within a safe operating voltage range (the rated voltage of this energy storage system is 48 VDC).
Around 6:00 AM, a while after sunrise, the PV systems charge the ESS, and the current display shows a positive value as shown in Fig. 8(b), indicating charging (a negative current indicates discharging). As the sunlight gradually intensifies, the charging current increases, reaching a maximum of 46 ADC around 9:15 AM. However, after 9:40 AM, because the energy storage system is fully charged, the system stops charging, and the current can be seen to drop instantly to 0 ADC. At this point, the cabins’ electricity is entirely supplied by solar energy.
@@ VII-A PV system replacement
    The existing PV modules of low efficiency (two channels, totaling 5.3 kWp) must be dismantled and removed. Existing usable brackets and fasteners should be collected and retained.
@@
    New PV arrays must have a capacity of 20 kWp or higher, and the PV modules must be Class-A with a module efficiency of 20% or higher. Product verification conforming to IEC performance and safety standards must be obtained.
@@
    Solar DC power protection devices such as surge absorbers, lightning bypass, short-circuit protectors, and overload circuit breakers must be installed in the solar DC power control enclosure.
@@
    Existing aluminum PV brackets must be assessed for any corrosion, paint peeling, cracking, deformation, or other conditions affecting structural strength, and a design should be implemented to maximize their retention and reuse to achieve environmental protection goals.
@@
    Outdoor connections to PV power lines must have sufficient high-voltage insulation and waterproofing capabilities. For example, the Stäubli MC4 solar-specific waterproof connector can be used, which has 1000 V high voltage and IP68 waterproofing capabilities.
@@
    To withstand the extremely harsh climate of high-altitude, low-temperature environments, the silicone material used for filling gaps must possess excellent weather resistance, maintaining its elasticity for extended periods. It must be ozone-resistant, ultra-violet-resistant (UV-resistant), and free from aging, cracking, or damage. It must be moisture-resistant and corrosion-resistant, preventing severe electrolytic corrosion. It must withstand temperature variations between freezing and boiling points while retaining the properties of the raw materials. It must have good adhesion and remain bonded for a long time without detaching.
@@ VII-B ESS replacement and enlargement
    The ESS cabinet can adopt a modular design to distribute the cabinet weight, facilitating collaborative transport uphill.
@@
    The battery can adopt a modular design to distribute the battery weight, facilitating collaborative transport uphill.
VII-C EMS delvelopment

### v2

@@ Abstract
Mountain huts are buildings located at high altitude, offering a place for hikers and providing shelter. Energy supply to mountain huts remains an ongoing issue. Using renewable energies could be an appropriate solution. Jiujiu Cabins, a famous mountain hut in Shei-Pa National Park, Taiwan, has operated an off-grid solar energy storage system (ESS) with lead-acid batteries. In 2021, a serious system failure took place, leading to no electricity. After a detailed on-site survey, a reorganization and repair project was implemented, and the energy system came back to operate normally. Meanwhile, an eco-friendly lithium iron phosphate battery (LFP battery) ESS replaces part of the lead-acid battery ESS, forming a hybrid ESS, making a better and greener off-grid solar ESS. In this case report, the energy architecture, detailed descriptions, and historical status of the system are provided. An on-site survey of the failed energy system, a system improvement project, and a future plan are listed.
@@
microgrid, renewable energy, photovoltaic system, energy storage system, hybrid energy storage system, lithium-ion battery, lithium iron phosphate battery, high mountain, mountain hut, high altitude, engineering.
@@ I Introduction
Mountain huts are buildings located at high altitude, in mountainous terrain, generally accessible only by foot, intended to provide food and shelter to mountaineers, hikers, and climbers [1]. They are known by many names, such as alpine hut, mountain shelter, mountain refuge, mountain lodge, mountain hostel, and mountain cabin. Mountain huts can provide some functionalities and services. The most important function is to provide shelter and simple sleeping berths for hikers to overcome the cold mountain environment at night. Many mountain huts are not staffed, particularly in remote areas. However, others have staff who provide hot meals, drinks, and other services, such as activity spaces (for holding lectures, etc.), clothing sales, and small items.
Energy supply in high mountains remains an open issue to be solved. Generally, grid connection is impractical, because the establishment and maintenance costs are too high. It also causes many environmental problems, e.g., the destruction of the natural ecology in the mountain. Diesel generators as energy suppliers have been used in high mountains for a long time despite their air pollution and noise problems. Due to the development of renewable energy (such as solar, wind, and hydropower), the usage of diesel generators is reduced, lowering the emissions of greenhouse gases (GHGs). However, owing to their fluctuating nature, most renewable energy sources exhibit intermittent features. To deal with the problem of renewable energy intermittency, energy storage systems (ESSs) are necessary [2, 3]. In the past, lead-acid batteries were heavily used as ESSs, accompanied by many environmental issues, e.g., poisoning, Pb leaks, contamination of the environment, and damage to the ecosystem [4, 5]. Recently, lithium-ion (Li-ion) batteries [6, 7], as a greener alternative, have started to replace lead-acid batteries in ESSs.
@@
In this manuscript, a brief introduction of the Jiujiu Cabins is given in Sec. II. In Sec. III, the current status of the energy architecture of the off-grid solar ESS is presented with a detailed description of each component. A historical development of energy systems is listed in Sec. IV. In 2021, Jiujiu Cabins experienced a brief period of power outage, and the investigation results of the system failure are listed in Sec. V. How to overhaul and improve the system, and the engineering work to restore power, is covered in Sec. VI. Possible future system upgrade plans are listed in Sec. VII. Summary and Outlook are given at last (Sec. VIII).
@@ II-A Current status
Jiujiu Cabins (coordinate: $24^{\circ}28^{\prime}08.7^{\prime\prime}$N, $121^{\circ}12^{\prime}29.9^{\prime\prime}$E) is an accommodation mountain hut in Taiwan, located in Shei-Pa National Park, in Tai’an Township, Miaoli County, at an altitude of 2,699 m (8,855 ft). It is named after the number 99 in its altitude [8]. The cabin complex belongs to the Hsinchu Branch, Forestry and Nature Conservation Agency, Ministry of Agriculture, Taiwan. It includes Chengkung Yurts, a restaurant (reception room), Longmen Inns, a machine room, toilet, etc., all of which are separate buildings (as shown in Fig. 2). Jiujiu Cabins uses a photovoltaic system and is equipped with a lithium-ion battery ESS to cope with the intermittent nature of renewable energy [8]. It can currently accommodate 150 people [9, 10]. It has clean running water (but currently no hot water provided) and a mobile phone communication point [11].
@@
Figure 2: Aerial photography of Jiujiu Cabins (2021). The cabin complex. It includes Chengkung Yurts, a restaurant (reception room), Longmen Inns, a machine room, toilet, etc., all of which are separate buildings. (Photographer: Yu-Chun Lin)
@@ III Architecture of current energy system
Due to the difficulty of power grid connection in high-altitude areas, energy supply remains a challenging problem. Off-grid solar ESSs [18, 19] provide an appropriate solution for using renewable energy in high-altitude areas and have been proven [18].
In order to restore the operation of the energy system, the Hsinchu Forest District Office, Forestry Bureau (renamed as the Forestry and Nature Conservation Agency in 2023) commissioned Super Double Power Technology Co., Ltd. to improve the existing solar energy system of Jiujiu Cabins in 2021 [20]. After restructuring and improvement, the off-grid solar energy storage system of Jiujiu Cabins is composed of two independent power supply systems combined with the same energy storage system architecture, as shown in Fig. 3. Among them, the first PV array inputs solar energy to hybrid solar inverters, which can transmit power to alternating current (AC) loads. The second PV array can transmit power to another set of AC loads through another hybrid solar inverter. The two sets of hybrid solar inverters share the same hybrid ESS, which consists of a newly established LFP battery [21] ESS and an existing lead-acid battery ESS. In addition, a new backup power input terminal is added, which can be connected to an external generator as a backup power source. The details of the energy architecture are described below.
@@ III-A Photovoltaic systems (PV systems)
In Jiujiu Cabins, there are two PV arrays, providing a total power generation capacity of 5.3 kWp. The first channel of the PV array is a ground-mounted PV system with a power generation capacity of 2.1 kWp. It is composed of 28 solar modules, each with a capacity of 75 Wp (SIEMENS Solar module SP75), as shown in Fig. 4. The second channel of the PV array is a ground-mounted PV system with a power generation capacity of 3.2 kWp. It is composed of 40 solar modules, each with a capacity of 80 Wp (SHARP Solar Module NE-80E2E), as shown in Fig. 5.
Figure 4: Ground-mounted PV system 1. PV array 1 is a ground-mounted PV system with a power generation capacity of 2.1 kWp. It is composed of 28 solar modules, each with a capacity of 75 Wp.
Figure 5: Ground-mounted PV system 2. PV array 2 is a ground-mounted PV system with a power generation capacity of 3.2 kWp. It is composed of 40 solar modules, each with a capacity of 80 Wp.
@@ III-F Backup power system
A diesel generator could be the backup power system. When the power from the PV system is insufficient (e.g., rainy days or cloudy days), the diesel generator can supply the insufficient power to the cabins. When the system encounters failures (e.g., temporary shutdown of the hybrid solar inverter), the diesel generator can supply the full power to the cabins.
@@ IV Historical development of energy systems
The Jiujiu Cabins had two existing electrical systems, designated as Old System 1 and Old System 2. These systems were likely installed by different manufacturers at different times. Owing to a malfunction in the Old System 1, power was unavailable. Another manufacturer then installed Old System 2 to supply power to the cabins’ 110 VAC loads. However, the inverter in Old System 2 also failed due to aging, causing power supply anomalies and no outputting electricity.
In order to restore the operation of the energy system, the Hsinchu Forest District Office commissioned Super Double Power Technology Co., Ltd. to reorganize, repair, and improve the existing solar energy system of Jiujiu Cabins in 2021 [20]. After the improvement project, it’s the current energy architecture (which is described in Sec. III in detail).
@@
Before Total: 5.3 kWp (PV) 15 kWh (Lead-acid) 3 kW No
@@ V On-site survey of energy systems
The Old System 1 can be considered as a microgrid system built on a 48 VDC DC bus. Solar panels (2.1 kWp) generate electricity, which is then charged via a solar charger (PWM) to a 48 VDC lead-acid battery ESS. A 48 VDC inverter converts the DC power to 110 VAC for output. Based on the property label, its installation date is Jan. 20, 2000.
The Old System 2 can be considered as a microgrid system built on a 24 VDC DC bus. After generating electricity from the solar panels (3.2 kWp), the electricity is charged to the 24 VDC lead-acid battery ESS through a solar charger (PWM). The DC power is then converted to 110 VAC AC power output by a 24 VDC inverter, with a rated power of 3 kW.
@@ VI-A Key issues
    The two existing PV arrays can be reused, but the series and parallel connections need to be adjusted within the inverter specifications.
@@ VI-B Main engineering subprojects
The solution involves restructuring the solar power lines, upgrading the inverters, restructuring the power supply lines at the cabins, and increasing the capacity of the ESS. However, this is not a simple matter of adding more batteries in parallel. It involves the parallel connection of old and new ESSs and the integration of the entire energy system.
The main engineering subprojects are: (1) adding LFP battery modules, (2) updating inverters, (3) modifying and reorganizing the existing PV array 1, (4) modifying and reorganizing the existing PV array 2, (5) changing control boxes (changing AC/DC lines), and (6) materials (hardware parts/connectors/DC lines). The following describes each subproject in detail:
@@
[…]
@@
[…]
@@
    Materials (hardware parts, connectors, and DC lines). (a) The old hardware parts, connectors, and DC lines were removed; (b) new hardware parts, connectors, DC lines were added according to the current decoration situation, including DC surge absorbers, DC dedicated leakage circuit breakers, coulomb meters, cloud EMS, and generator charging input terminals.
@@ VI-C EMS data analysis
In this project, due to budget constraints, the self-developed cloud-native EMS only records the total voltage, input/output current, and input/output power of the ESS. Because the cabins’ energy usage has a certain cyclical characteristic on a daily basis, a one-day data analysis is used to assess the system’s condition, as shown in Fig. 8.
@@
The 24-hour data shows that the highest total voltage of the ESS does not exceed 54 VDC (as shown in Fig. 8(a)), indicating that the BMS provides stable control over the LFP battery, keeping it within a safe operating voltage range (the rated voltage of this ESS is 48 VDC).
Around 6:00 AM, a while after sunrise, the PV systems charge the ESS, and the current display shows a positive value as shown in Fig. 8(b), indicating charging (a negative current indicates discharging). As the sunlight gradually intensifies, the charging current increases, reaching a maximum of 46 ADC around 9:15 AM. However, after 9:40 AM, because the ESS is fully charged, the system stops charging, and the current can be seen to drop instantly to 0 ADC. At this point, the cabins’ electricity is entirely supplied by solar energy.
@@ VII-A PV system replacement
    The existing PV modules of low efficiency (two channels, totaling 5.3 kWp) should be dismantled and removed. Existing usable brackets and fasteners should be collected and retained.
@@
    New PV arrays should have a capacity of 20 kWp or higher, and the PV modules should be Class-A with a module efficiency of 20% or higher. Product verification conforming to IEC performance and safety standards should be obtained.
@@
    Solar DC power protection devices such as surge absorbers, lightning bypass, short-circuit protectors, and overload circuit breakers should be installed in the solar DC power control enclosure.
@@
    Existing aluminum PV brackets should be assessed for any corrosion, paint peeling, cracking, deformation, or other conditions affecting structural strength, and a design should be implemented to maximize their retention and reuse to achieve environmental protection goals.
@@
    Outdoor connections to PV power lines should have sufficient high-voltage insulation and waterproofing capabilities. For example, the Stäubli MC4 solar-specific waterproof connector can be used, which has 1000 V high voltage and IP68 waterproofing capabilities.
@@
    To withstand the extremely harsh climate of high-altitude, low-temperature environments, the silicone material used for filling gaps should possess excellent weather resistance, maintaining its elasticity for extended periods. It should be ozone-resistant, ultra-violet-resistant (UV-resistant), and free from aging, cracking, or damage. It should be moisture-resistant and corrosion-resistant, preventing severe electrolytic corrosion. It should withstand temperature variations between freezing and boiling points while retaining the properties of the raw materials. It should have good adhesion and remain bonded for a long time without detaching.
@@ VII-B ESS replacement and enlargement
    The ESS cabinet can adopt a modular design to distribute the cabinet weight. This allows the mountain porters to successfully bring the materials up the mountain.
@@
    The battery can adopt a modular design to distribute the battery weight. This allows the mountain porters to successfully bring the materials up the mountain.
VII-C EMS development